OpenStudy (aravindg):

Show that each one of the following is an even function: (i) f(x) = 1/x2 for all x Є R – {0} (ii) f(x)=ІxІ/(x2 + 1) for all x Є R

OpenStudy (zarkon):

show f(-x)=f(x)

OpenStudy (amistre64):

even function retain the same value when you negate the x f(-x) = f(x)

OpenStudy (aravindg):

how ????

OpenStudy (aravindg):

show please

OpenStudy (zarkon):

\[f(-x)=\frac{1}{(-x)^2}=\frac{1}{x^2}=f(x)\]

OpenStudy (amistre64):

x = a ; and x = -a; solve for f(a) and f(-a) perhaps

OpenStudy (zarkon):

\[f(-x)=\frac{|-x|}{(-x)^2+1}=\cdots\]